OpenStudy (anonymous):

Help? I cant figure out the correct answer Find an equation for the nth term of the arithmetic sequence. a_19 = -58, a_21 = -164

OpenStudy (anonymous):

and these are the options: an = 896 - 53(n - 2) an = 896 - 53(n - 1) an = 896 + 53(n + 1) an = 896 - 53(n + 1)

OpenStudy (er.mohd.amir):

a_19 means 19 th term and so on

OpenStudy (er.mohd.amir):

but n th term is =a+(n-1)d where a is first term and d is common diff.

OpenStudy (tkhunny):

Walk yourself through it. \(a_{1} = a\) \(a_{2} = a + d\) \(a_{3} = a + d + d = a + 2d\) \(a_{4} = a + 3d = a + (4-1)d\) \(a_{5} = a + 4d = a + (5-1)d\) ... \(a_{n} = a + (n-1)d\) ... \(a_{19} = a + (19-1)d = a + 18d = -58\) \(a_{20} = a + (20-1)d = a + 19d\) \(a_{21} = a + (21-1)d = a + 20d = -164\) That is enough information to find 'a' and 'd'.